# Griesmer and Optimal Linear Codes from the Affine Solomon-Stiffler Construction

### v1

@@ 1.2 Related works
Several infinite families of $q$-ary linear Griesmer codes or optimal codes have been constructed in [26, 27, 25, 18, 23] as Gray images of two-weight codes over finite rings. Weight distributions of these Griesmer and optimal codes were determined in these two papers [26, 18]. In [26], Shi, Guan and Solé constructed optimal codes in [26, Theorem 6.1, 6.3] via Gray images of some two-weight codes over the finite ring ${\bf F}_{p}+u{\bf F}_{p}$ and determined their exact weight distributions. The main contribution of [18] is the construction of Griemser codes in [18, Theorem 3.5], based a similar method. Similarly, Mesnager, Qian and Cao constructed optimal codes in [23, Theorem 5.4] via Gray images of codes over a finite chain ring. In [15], more general trace codes, which reach the Griesmer bound or are distance-optimal were constructed. In [25, Theorm 2] and [24, Proposition 4.3, 5)], some binary $[2^{k}-2^{u},k-1,2^{k-1}-2^{u-1}]_{2}$ codes were constructed.
@@ 1.3 Contributions and the organization of the paper
2) In Sections 2 and 3, we show that these Griesmer codes and optimal codes constructed in [26, 18] and very special case codes in our construction have the same parameters and weight distributions. Optimal linear $q$-ary codes constructed in [23] have same parameters as very special codes in our construction. It is also indicated that Griesmer and optimal codes constructed in this paper are less restrictive than that constructed in [15]. In Section 7, we construct much more general distance-optimal binary codes than these constructed in [25, 24]. Our distance-optimal codes are codimension one or two subcodes of binary Solomon-Stiffler codes.
@@
The paper is organized as follows. In Section 2, affine Solomon-Stiffler codes are introduced. Though these codes are not Griesmer as projective Solomon-Stiffler codes, their Griesmer defects can be upper bounded in some cases. In Section 3, several infinite families of optimal affine Solomon-Stiffler codes are given and their exact weight distributions are determined. In Section 4, modified affine Solomon-Stiffler codes are presented. Comparing to affine Solomon-Stiffler codes, their Griesmer defects decrease. Therefore, modified Solomon-Stiffler codes are more closer to the Griesmer bound. In Section 5, some infinite families of Griesmer modified affine Solomon-Stiffler codes are constructed and their weight distributions are determined. In Section 6, an infinite family of almost optimal codes is given, In Section 7, distance-optimal binary codes are constructed from codimension one subcodes of Griesmer codes. Section 8 concludes the paper.
@@ 2 Affine Solomon-Stiffler codes
Proof. For any hyperplane $H\subset{\bf F}_{q}^{k}$, it is clear that $|H\bigcap S|$ is minimal if and only if $H\bigcap S_{i}$ is a hyperplane in $S_{i}$. Then $H-H\bigcap S$ is maximal, when $H\bigcap S_{i}$ is a hyperplane in $S_{i}$, for $i=1,\ldots,h$. The weight of the nonzero codeword corresponding to $H$ is $|S|-|(H-H\bigcap S)|$. The first conclusion follows directly. Since $|H\bigcap S|$ has at most $2^{h}$ possible values, the second conclusion is proved.
@@
Lemma 2 Let ${\bf C}$ be a linear $[n,k,d]_{q}$ code with the Griesmer defect $g$. Then the optimal distance of a linear $[n,k]_{q}$ code is at most $d+g+1$.
@@
Proof. We calculate the Griesmer defect. Notice that $\lceil\frac{(q-1)(q^{k-1}-\Sigma_{i=s_{1}+1}^{s+1+s_{2}}q^{u_{i}-1})}{q^{i}}\rceil\leq(q-1)\lceil\frac{(q^{k-1}-\Sigma_{i=s_{1}+1}^{s_{1}+s_{2}}q^{u_{i}-1})}{q^{i}}\rceil-(s_{2}-1)$, only when $i=u_{s_{1}+1}$. On the other hand, if $u_{s_{1}+s_{2}+\cdots+s_{j}+1}\geq u_{s_{1}+s_{2}+\cdots+s_{j}}+2$, this kind of decreasing happens only when $i=u_{s_{j}+1}$. Then the conclusion follows.
@@
Let $q$ be a prime power and $s$ be a positive integer satisfying $s\geq 3$. The affine Solomon-Stiffler code associated with parameters $(2s-2,s)$ is a linear Griesmer $[q^{2s-2}-q^{s-1},2(s-1),(q-1)(q^{2s-3}-q^{s-2}]_{q}$ code. The $q$-copy repetition code is a linear $[q^{2s-1}-q^{s},2(s-1),(q-1)(q^{2s-2}-q^{s-1})]_{q}$ code. This code is optimal when $s\geq q$ as argued in [23, Theorem 5.4]. Comparing these two codes, they have the same parameters. The construction of optimal codes in [23] is based on Griesmer codes over a finite chain ring. Our construction is simpler.
In [15, Theorem 1], similar codes with some strong restrictions on dimensions of subfields of ${\bf F}_{q^{k}}$ are studied. Only codes in the case when $u_{i}$ is not a factor of $u_{j}$, if $j>i$, or all $u_{i}$’s are the same, are studied. Comparing to [15], we need the key condition 1). However it is clear that Theorem 2 gives more general Griesmer and optimal codes than trace Griesmer and optimal codes given in [15].
@@ 3 Optimal few-weight affine Solomon-Stiffler codes
Let us consider the case $k=2m$, $u_{1}=u_{2}=m$. Then we construct a linear $[q^{2m}-2q^{m}+1,2m,(q-1)(q^{2m-1}-2q^{m-1})]_{q}$ code. From Theorem 2, the Griesmer defect is one and this is an optimal code. These codes have the same parameters as optimal codes constructed by Shi, Guan and Solé in [26, Theorem 6.1]. The optimal codes in [26, Theorem 6.2], are just a repetition copy of the above code. Their $q$ needs to be an odd prime satisfying $q\equiv 3$ $od$ $4$, and $m$ needs to be odd. It should be indicated that our codes are more general. Moreover optimal codes in [26] were constructed via Gray images of two-weight codes over a finite ring.
@@ 00 $1$
Notice that the weight distribution of this infinite family of optimal two-weight $q$-ary linear codes is the same as the weight distribution Table II of optimal two-weight codes constructed in [26, Theorem 6.1]. Their construction via Gray images of two-weight codes over the finite ring ${\bf F}_{p}+u{\bf F}_{p}$, seems unnecessary, since their strong conditions on $q$ and $m$ are not essential. When $m=2r$, codes constructed in [15, Proposition 5] are these codes in Theorem 3.
@@
Example 6 When $q=3$ and $m=2$, we construct an optimal ternary three-weight $[208,5,138]_{3}$ code, see [9]. This is a Griesmer code. Three weights are $138,144,154$, $A_{138}=242$, $A_{144}=26$ and $A_{154}=8$. The optimal $[208,5,138]_{3}$ code in [9] is from a stored generator matrix.
@@
Theorem 5 1) Let $q>2$ be a prime power. Let $m$ be a positive integer. Let $(k,{\bf u})=(3m+1,m,m,m+1)$. Then we construct an infinite family of optimal five-weight $[q^{3m+1}-2q^{m}-q^{m+1}+2,3m+1,(q-1)(q^{3m-1}-2q^{m-1}-q^{m})]_{q}$ codes. The weight distribution is in the following table.
@@ 4 Modified affine Solomon-Stiffler codes
Notice that this is the same as the Griesmer codes constructed in [18, Theorem 3.5]. The weight distribution is the same as that in [18, Theorem 3.3]. The construction in [18], as that in [26], is very technique, and some restrictions on $e$ and $m$ were needed.
Example 10 When $q=5$, $(k,{\bf u})=(4,1,1,1,1)$, we construct a linear $[152,4,121]_{5}$ code. This code is Griesmer code and non-Solomon-Stiffler.
@@ 5 Few-weight Griesmer codes from the modified affine Solomon-Stiffler construction
In this paper, we construct some two-weight and three-weight Griesmer codes and determine their weight distributions.
@@
Corollary 4 Let $q$ be a prime power and $e$ be a divisor of $q-1$. Then we construct an infinite family of Griesmer $[\frac{e(q^{2m}-2q^{m}+1)}{q-1},2m,e(q^{2m-1}-q^{m-1})]_{q}$ codes. The weight distribution is in the following table.
@@ 6 Modified affine Solomon-Stiffler codes when $u_{1}=u_{2}\cdots=u_{h}=1$, $h$ large
In this section, we consider affine Solomon-Stiffler codes associated with $u_{1}=\cdots=u_{h}=1$, $h$ is large positive integer satisfying $h<q^{k-1}$. We prove the following result.
@@
Proof. From the proof of Theorem 1, we have $d\geq q^{k-1}-2q$. If the maximal possible distance of a linear $[\frac{q^{k}-1}{q-1}-2q,k]_{q}$ is $q^{k-1}-2q+2$, then from the Griesmer bound, $\frac{q^{k}-1}{q-1}-2q\geq q^{k-1}-2q+2+q^{k-2}-1+q^{k-3}+\cdots+q+1=\frac{q^{k}-1}{q-1}-2q+1$. This is a contradiction.
@@
$\Sigma_{i=0}^{k-1}\lfloor\frac{2^{k-1}-2}{2^{i}}\rfloor=2^{k}-4>2^{k}-5.$
@@
When $q=3,h=6$, we construct an infinite family of linear $[\frac{3^{k}-1}{2},k,3^{k-1}-6]_{3}$ code with the Griesmer defect at most $2$. This code is almost optimal, since
$\Sigma_{i=0}^{k-1}\lfloor\frac{3^{k-1}-4}{3^{i}}\rfloor=\frac{3^{k}-1}{2}-5>\frac{3^{k}-1}{2}-6.$
@@
7 Distance-optimal codes as codimension one subcodes of Griesmer codes
@@
Theorem 9 Let ${\bf C}$ be a binary linear Griesmer $[n,k,2d]_{2}$ code satisfying $k\geq 3$ and $0<d\leq 2^{k-2}$. Then any dimension $k-1$ subcode of ${\bf C}$ has minimum weight $2d$ and is distance-optimal.
@@
$n\geq\Sigma_{i=0}^{k-2}\lceil\frac{2d+1}{2^{i}}\rfloor\geq 2+\Sigma_{i=0}^{k-2}\lfloor\frac{2d}{2^{i}}\rfloor=1+\Sigma_{i=0}^{k-1}\lfloor\frac{2d}{2^{i}}\rfloor.$
@@
$n=\Sigma_{i=0}^{k-1}\lfloor\frac{2d}{2^{i}}\rfloor,$
@@
Distance-optimal codes constructed in [25, Theorem 2] and [24, Proposition 5.3, 5)] have the same parameters as above codes. Their constructions are complicated and technique. Actually, more general such distance-optimal binary codes are constructed in the following result.
Corollary 8 Let ${\bf C}$ be a binary Solomon-Stiffler $[2^{k}-1-\Sigma_{i=1}^{h}(2^{u_{i}}-1),k,2^{k-1}-\Sigma_{i=1}^{h}2^{u_{i}-1}]_{2}$ code constructed in [28]. We assume that $k\geq 3$, $2\leq u_{1}<u_{2}<\cdots<u_{h}<k$, Then we construct many dimension $k-1$ distance-optimal binary $[2^{k}-1-\Sigma_{i=1}^{h}(2^{u_{i}}-1),k-1,2^{k-1}-\Sigma_{i=1}^{h}2^{u_{i}-1}]_{2}$ code.
@@
Theorem 10 Let ${\bf C}$ be a binary linear Griesmer $[n,k,4d]_{2}$ code satisfying $k\geq 4$ and $0<d\leq 2^{k-4}$. Then any dimension $k-2$ subcode of ${\bf C}$ has minimum weight $4d$ and is distance-optimal.
@@
$n\geq\Sigma_{i=0}^{k-3}\lceil\frac{4d+1}{2^{i}}\rfloor\geq 3+\Sigma_{i=0}^{k-3}\lfloor\frac{4d}{2^{i}}\rfloor=1+\Sigma_{i=0}^{k-1}\lfloor\frac{4d}{2^{i}}\rfloor.$
@@
$n=\Sigma_{i=0}^{k-1}\lfloor\frac{4d}{2^{i}}\rfloor,$
@@
Corollary 9 Let ${\bf C}$ be a binary Solomon-Stiffler $[2^{k}-1-\Sigma_{i=1}^{h}(2^{u_{i}}-1),k,2^{k-1}-\Sigma_{i=1}^{h}2^{u_{i}-1}]_{2}$ code constructed in [28]. We assume that $k\geq 4$, $u_{h}=k-1$ and $23\leq u_{1}<u_{2}<\cdots<u_{h-1}\leq k-$ 2, Then we construct many dimension $k-1$ distance-optimal binary $[2^{k}-1-\Sigma_{i=1}^{h}(2^{u_{i}}-1),k-1,2^{k-1}-\Sigma_{i=1}^{h}2^{u_{i}-1}]_{2}$ code.
@@
Theorem 11 1) Let ${\bf C}$ be a $q$-ary linear Griesmer $[n,k,qd]_{q}$ code satisfying $k\geq 3$ and $0<qd\leq q^{k-1}$. Then any dimension $k-1$ subcode of ${\bf C}$ has minimum weight $qd$ and is distance-optimal.
2) Let ${\bf C}$ be a $q$-ary linear Griesmer $[n,k,q^{2}d]_{q}$ code satisfying $k\geq 4$ and $0<q^{2}d\leq q^{k-2}$. Then any dimension $k-2$ subcode of ${\bf C}$ has minimum weight $q^{2}d$ and is distance-optimal.
From $q$-ary projective Solomon-Stiffler codes satisfying certain restrictions, many distance-optimal $q$-ary linear codes are constructed.

### v2

@@ 1.2 Related works
Several infinite families of $q$-ary linear Griesmer codes or optimal codes have been constructed in [27, 28, 26, 19, 24] as Gray images of two-weight codes over finite rings. Weight distributions of these Griesmer and optimal codes were determined in these two papers [27, 19]. In [27], Shi, Guan and Solé constructed optimal codes in [27, Theorem 6.1, 6.3] via Gray images of some two-weight codes over the finite ring ${\bf F}_{p}+u{\bf F}_{p}$ and determined their exact weight distributions. The main contribution of [19] is the construction of Griemser codes in [19, Theorem 3.5], based a similar method. Similarly, Mesnager, Qian and Cao constructed optimal codes in [24, Theorem 5.4] via Gray images of codes over a finite chain ring. In [16], more general trace codes, which reach the Griesmer bound or are distance-optimal were constructed. In [26, Theorm 2] and [25, Proposition 4.3, 5)], some binary $[2^{k}-2^{u},k-1,2^{k-1}-2^{u-1}]_{2}$ codes were constructed. In a recent paper [15], punctured and shortening optimal codes were obtained from Griesmer cyclic codes.
@@ 1.3 Contributions and the organization of the paper
2) In Sections 2 and 3, we show that these Griesmer codes and optimal codes constructed in [27, 19] and very special case codes in our construction have the same parameters and weight distributions. Optimal linear $q$-ary codes constructed in [24] have same parameters as very special codes in our construction. It is also indicated that Griesmer and optimal codes constructed in this paper are less restrictive than that constructed in [16]. In Section 7, we construct much more general distance-optimal binary codes than these constructed in [26, 25]. Our distance-optimal codes are codimension one or two subcodes of binary Solomon-Stiffler codes. Distance-optimal punctured codes of Griesmer codes are also discussed in Section 8. Our results in Section 7 and 8 are more general than main results of [15].
@@
The paper is organized as follows. In Section 2, affine Solomon-Stiffler codes are introduced. Though these codes are not Griesmer as projective Solomon-Stiffler codes, their Griesmer defects can be upper bounded in some cases. In Section 3, several infinite families of optimal affine Solomon-Stiffler codes are given and their exact weight distributions are determined. In Section 4, modified affine Solomon-Stiffler codes are presented. Comparing to affine Solomon-Stiffler codes, their Griesmer defects decrease. Therefore, modified Solomon-Stiffler codes are more closer to the Griesmer bound. In Section 5, some infinite families of Griesmer modified affine Solomon-Stiffler codes are constructed and their weight distributions are determined. In Section 6, an infinite family of almost optimal codes is given, In Section 7, distance-optimal binary codes are constructed from codimension one subcodes of Griesmer codes. In Section 8, distance-optimal punctured codes of Griesmer codes are give. Section 9 concludes the paper.
@@ 2 Affine Solomon-Stiffler codes
Proof. For any hyperplane $H\subset{\bf F}_{q}^{k}$, it is clear that $|H\bigcap S_{i}|$ is minimal if and only if $H\bigcap S_{i}$ is a hyperplane in $S_{i}$. Then $H-H\bigcap S$ is maximal, when each $H\bigcap S_{i}$ is a hyperplane in $S_{i}$, for $i=1,\ldots,h$. The weight of the nonzero codeword corresponding to $H$ is $|F_{2}^{k}-S|-|(H-H\bigcap S)|$. The first conclusion follows directly. Since each $|H\bigcap S_{i}|$ has two possible values, then $|H\bigcap S|$ has at most $2^{h}$ possible values. The second conclusion is proved.
@@
Lemma 2 Let ${\bf C}$ be a linear $[n,k,d]_{q}$ code with the Griesmer defect $g$. Then the optimal distance of a linear $[n,k]_{q}$ code is at most $d+g$.
@@
Proof. We calculate the Griesmer defect. Notice that $\lceil\frac{(q-1)(q^{k-1}-\Sigma_{i=s_{1}+1}^{s_{1}+s_{2}}q^{u_{i}-1})}{q^{i}}\rceil\leq(q-1)\lceil\frac{(q^{k-1}-\Sigma_{i=s_{1}+1}^{s_{1}+s_{2}}q^{u_{i}-1})}{q^{i}}\rceil-(s_{2}-1)$, only when $i=u_{s_{1}+1}$. On the other hand, if $u_{s_{1}+s_{2}+\cdots+s_{j}+1}\geq u_{s_{1}+s_{2}+\cdots+s_{j}}+2$, this kind of decreasing happens only when $i=u_{s_{j}+1}$. Then the conclusion follows.
@@
Let $q$ be a prime power and $s$ be a positive integer satisfying $s\geq 3$. The affine Solomon-Stiffler code associated with parameters $(2s-2,s)$ is a linear Griesmer $[q^{2s-2}-q^{s-1},2(s-1),(q-1)(q^{2s-3}-q^{s-2}]_{q}$ code. The $q$-copy repetition code is a linear $[q^{2s-1}-q^{s},2(s-1),(q-1)(q^{2s-2}-q^{s-1})]_{q}$ code. This code is optimal when $s\geq q$ as argued in [24, Theorem 5.4]. Comparing these two codes, they have the same parameters. The construction of optimal codes in [24] is based on Griesmer codes over a finite chain ring. Our construction is simpler.
In [16, Theorem 1], similar codes with some strong restrictions on dimensions of subfields of ${\bf F}_{q^{k}}$ are studied. Only codes in the case when $u_{i}$ is not a factor of $u_{j}$, if $j>i$, or all $u_{i}$’s are the same, are studied. Comparing to [16], we need the key condition 1). However it is clear that Theorem 2 gives more general Griesmer and optimal codes than trace Griesmer and optimal codes given in [16].
@@ 3 Optimal few-weight affine Solomon-Stiffler codes
Let us consider the case $k=2m$, $u_{1}=u_{2}=m$. Then we construct a linear $[q^{2m}-2q^{m}+1,2m,(q-1)(q^{2m-1}-2q^{m-1})]_{q}$ code. From Theorem 2, the Griesmer defect is one and this is an optimal code. These codes have the same parameters as optimal codes constructed by Shi, Guan and Solé in [27, Theorem 6.1]. The optimal codes in [27, Theorem 6.2], are just a repetition copy of the above code. Their $q$ needs to be an odd prime satisfying $q\equiv 3$ $od$ $4$, and $m$ needs to be odd. It should be indicated that our codes are more general. Moreover optimal codes in [27] were constructed via Gray images of two-weight codes over a finite ring.
@@ 00 $1$
Notice that the weight distribution of this infinite family of optimal two-weight $q$-ary linear codes is the same as the weight distribution Table II of optimal two-weight codes constructed in [27, Theorem 6.1]. Their construction via Gray images of two-weight codes over the finite ring ${\bf F}_{p}+u{\bf F}_{p}$, seems unnecessary, since their strong conditions on $q$ and $m$ are not essential. When $m=2r$, codes constructed in [16, Proposition 5] are these codes in Theorem 3.
@@
Example 6 When $q=3$ and $m=2$, we construct an optimal ternary three-weight $[208,5,138]_{3}$ code, see [9]. This is a Griesmer code. Three weights are $138,144,156$, $A_{138}=208$, $A_{144}=26$ and $A_{156}=8$. The optimal $[208,5,138]_{3}$ code in [9] is from a stored generator matrix.
@@
Theorem 5 1) Let $q>2$ be a prime power. Let $m$ be a positive integer. Let $(k,{\bf u})=(3m+1,m,m,m+1)$. Then we construct an infinite family of optimal five-weight $[q^{3m+1}-2q^{m}-q^{m+1}+2,3m+1,(q-1)(q^{3m}-2q^{m-1}-q^{m})]_{q}$ codes. The weight distribution is in the following table.
@@ 4 Modified affine Solomon-Stiffler codes
Notice that this is the same as the Griesmer codes constructed in [19, Theorem 3.5]. The weight distribution is the same as that in [19, Theorem 3.3]. The construction in [19], as that in [27], is very technique, and some restrictions on $e$ and $m$ were needed.
Example 10 When $q=5$, $(k,{\bf u})=(4,1,1,1,1),e=1$, we construct a linear $[152,4,121]_{5}$ code. This code is Griesmer code and non-Solomon-Stiffler.
@@ 5 Few-weight Griesmer codes from the modified affine Solomon-Stiffler construction
In this section, we construct some two-weight and three-weight Griesmer codes and determine their weight distributions.
@@
Corollary 4 Let $q$ be a prime power and $e$ be a divisor of $q-1$. Then we construct an infinite family of Griesmer $[\frac{e(q^{2m}-2q^{m}+1)}{q-1},2m,e(q^{2m-1}-2q^{m-1})]_{q}$ codes. The weight distribution is in the following table.
@@ 6 Modified affine Solomon-Stiffler codes when $u_{1}=u_{2}\cdots=u_{h}=1$, $h$ large
In this section, we consider affine Solomon-Stiffler codes associated with $u_{1}=\cdots=u_{h}=1$, $h$ is a large positive integer satisfying $h<q^{k-1}$. We prove the following result.
@@
Proof. From the proof of Theorem 1, we have $d\geq q^{k-1}-2q$. If the maximal possible distance of a linear $[\frac{q^{k}-1}{q-1}-2q,k]_{q}$ is $q^{k-1}-2q+2$, then from the Griesmer bound, $\frac{q^{k}-1}{q-1}-2q\leq q^{k-1}-2q+2+q^{k-2}-1+q^{k-3}+\cdots+q+1=\frac{q^{k}-1}{q-1}-2q+1$. This is a contradiction.
@@
$\Sigma_{i=0}^{k-1}\lceil\frac{2^{k-1}-2}{2^{i}}\rceil=2^{k}-4>2^{k}-5.$
@@
When $q=3,h=6$, we construct an infinite family of linear $[\frac{3^{k}-1}{2}-6,k,3^{k-1}-6]_{3}$ code with the Griesmer defect at most $2$. This code is almost optimal, since
$\Sigma_{i=0}^{k-1}\lceil\frac{3^{k-1}-4}{3^{i}}\rceil=\frac{3^{k}-1}{2}-5>\frac{3^{k}-1}{2}-6.$
@@
7 Distance-optimal codes as codimension one or two subcodes in Griesmer codes
@@
Theorem 9 Let ${\bf C}$ be a binary linear Griesmer $[n,k,2d]_{2}$ code satisfying $k\geq 3$ and $d\leq 2^{k-2}$. Then any dimension $k-1$ subcode of ${\bf C}$ has minimum weight $2d$ and is distance-optimal. The Griesmer defect of this shortening code is $1$.
@@
$n\geq\Sigma_{i=0}^{k-2}\lceil\frac{2d+1}{2^{i}}\rceil\geq 2+\Sigma_{i=0}^{k-2}\lceil\frac{2d}{2^{i}}\rceil=1+\Sigma_{i=0}^{k-1}\lceil\frac{2d}{2^{i}}\rceil.$
@@
$n=\Sigma_{i=0}^{k-1}\lceil\frac{2d}{2^{i}}\rceil,$
@@
Distance-optimal codes constructed in [26, Theorem 2] and [25, Proposition 5.3, 5)] have the same parameters as above codes. Their constructions are complicated and technique. Actually, more general such distance-optimal binary codes are constructed in the following result. By the way, optimal binary three-weight $[3\cdot 2^{3m-1},3m,3\cdot 2^{3m-2}]_{2}$ codes constructed in [29] have the same parameters as repetition codes of the 1st order RM $[2^{3m-1},3m,2^{3m-2}]_{2}$ codes.
Corollary 8 Let ${\bf C}$ be a binary Solomon-Stiffler $[2^{k}-1-\Sigma_{i=1}^{h}(2^{u_{i}}-1),k,2^{k-1}-\Sigma_{i=1}^{h}2^{u_{i}-1}]_{2}$ code constructed in [30]. We assume that $k\geq 3$, $2\leq u_{1}<u_{2}<\cdots<u_{h}<k$, Then we construct many dimension $k-1$ distance-optimal binary $[2^{k}-1-\Sigma_{i=1}^{h}(2^{u_{i}}-1),k-1,2^{k-1}-\Sigma_{i=1}^{h}2^{u_{i}-1}]_{2}$ codes. The Griesmer defect is one.
@@
Theorem 10 Let ${\bf C}$ be a binary linear Griesmer $[n,k,4d]_{2}$ code satisfying $k\geq 4$ and $d\leq 2^{k-4}$. Then any dimension $k-2$ subcode of ${\bf C}$ has minimum weight $4d$ and is distance-optimal. The Griesmer defect of $k-2$ dimension subcode is two.
@@
$n\geq\Sigma_{i=0}^{k-3}\lceil\frac{4d+1}{2^{i}}\rceil\geq 3+\Sigma_{i=0}^{k-3}\lceil\frac{4d}{2^{i}}\rceil=1+\Sigma_{i=0}^{k-1}\lceil\frac{4d}{2^{i}}\rceil.$
@@
$n=\Sigma_{i=0}^{k-1}\lceil\frac{4d}{2^{i}}\rceil,$
@@
Corollary 9 Let ${\bf C}$ be a binary Solomon-Stiffler $[2^{k}-1-\Sigma_{i=1}^{h}(2^{u_{i}}-1),k,2^{k-1}-\Sigma_{i=1}^{h}2^{u_{i}-1}]_{2}$ code constructed in [30]. We assume that $k\geq 4$, $u_{h}=k-1$ and $23\leq u_{1}<u_{2}<\cdots<u_{h-1}\leq k-$ 2, Then we construct many dimension $k-2$ distance-optimal binary $[2^{k}-1-\Sigma_{i=1}^{h}(2^{u_{i}}-1),k-1,2^{k-1}-\Sigma_{i=1}^{h}2^{u_{i}-1}]_{2}$ codes. The Griesmer defect of the dimension $k-2$ subcode is two.
@@
Theorem 11 1) Let ${\bf C}$ be a $q$-ary linear Griesmer $[n,k,qd]_{q}$ code satisfying $k\geq 3$ and $d\leq q^{k-2}$. Then any dimension $k-1$ subcode of ${\bf C}$ has minimum weight $qd$ and is distance-optimal. The Griesmer defect of this codimension one subcode is one.
2) Let ${\bf C}$ be a $q$-ary linear Griesmer $[n,k,q^{2}d]_{q}$ code satisfying $k\geq 4$ and $d\leq q^{k-4}$. Then any dimension $k-2$ subcode of ${\bf C}$ has minimum weight $q^{2}d$ and is distance-optimal. The Griesmer defect of this codimension two subcode is two
From $q$-ary projective Solomon-Stiffler codes satisfying certain restrictions, many distance-optimal $q$-ary linear codes are constructed. All these distance-optimal codes have positive Griesmer defects.
Results in this section shows that, when some optimal codes with new parameters were constructed, it would be better to check if these distance-optimal codes (with respect to the Griesmer bound) are small codimension subcodes of known Griesmer codes.
8 Distance-optimal codes as punctured codes of Griesmer codes
In this section, we prove the following result for punctured codes of $q$-ary Griesmer codes.
Theorem 12 1) Let ${\bf C}$ be a $q$-ary linear Griesmer $[n,k,qd]_{q}$ code satisfying $d=qd^{\prime}>1$. Then the punctured code at a position in the support of a minimum weight codeword of ${\bf C}$ is a linear $[n-1,k,qd^{\prime}-1]_{q}$ Griesmer code.
2) Let ${\bf C}$ be a $q$-ary linear Griesmer $[n,k,qd^{\prime}+1]_{q}$ code satisfying $k\geq 3$. Suppose that $d^{\prime}$ satisfies that $d^{\prime}\geq 1$, $d^{\prime}\equiv d^{\prime\prime}$ $mod$ $q$, where $d^{\prime\prime}<q-1$. Then the punctured code ${\bf C}$ at a position in the support of a minimum weight codeword of ${\bf C}$ has the Griesmer defect one. It is distance-optimal.
Proof. We have
$\Sigma_{i=0}^{k-1}\lceil\frac{qd^{\prime}-1}{q^{i}}\rceil=qd^{\prime}-1+\lceil\frac{qd^{\prime}-1}{q}\rceil+\cdots+\lceil\frac{qd^{\prime}-1}{q^{k-1}}\rceil=\Sigma_{i=0}^{k-1}\lceil\frac{qd^{\prime}}{q^{i}}-1=n-1.$
The conclusion 1) is proved.
Suppose that $d^{\prime}=qd_{1}+d^{\prime\prime}$, where $d^{\prime\prime}<q-1$. Then,
$\Sigma_{i=0}^{k-1}\lceil\frac{qd^{\prime}}{q^{i}}\rceil=qd^{\prime}+d^{\prime}+\lceil\frac{q^{2}d_{1}+qd^{\prime\prime}}{q^{2}}\rceil+\cdots+\lceil\frac{q^{2}d_{1}+qd^{\prime\prime}}{q^{k-1}}\rceil=\Sigma_{i=0}^{k-1}\lceil\frac{qd^{\prime}+1}{q^{i}}\rceil-2.$
The Griesmer defect of the punctured code is one.
If there is a linear $[n-1,k,qd^{\prime}+1]_{2}$ code, then it is a contradiction to the condition that the linear $[n,k,qd^{\prime}+1]_{q}$ code is Griesmer. Therefore the punctured $[n-1,k,qd^{\prime}]_{q}$ code is distance-optimal.
When Theorem 12 is applied to binary Solomon codes, then we have the following result.
Corollary 10 Let ${\bf C}$ be a binary Solomon-Stiffler $[2^{k}-1-\Sigma_{i=1}^{h}(2^{u_{i}}-1),k,2^{k-1}-\Sigma_{i=1}^{h}2^{u_{i}-1}]_{2}$ code constructed in [30]. We assume that $k\geq 3$, $u_{1}=1,u_{2}=2$ and $13\leq u_{3}<\cdots<u_{h-1}\leq k-$ 1, Then we construct many dimension $k$ distance-optimal binary $[2^{k}-1-\Sigma_{i=1}^{h}(2^{u_{i}}-1),k,2^{k-1}-\Sigma_{i=1}^{h}2^{u_{i}-1}-1]_{2}$ codes, as punctured codes. The Griesmer defect of these punctured codes is one.